OpenStudy (anonymous):

What is the solution to the equation? 5(x-4)-4x=18-23

OpenStudy (anonymous):

@DullJackel09

OpenStudy (anonymous):

multiply 5x-20-4x=18-23 sum equla terms, x=20+18-23 x=15 the solution is x=15